Calculation of Value-Added and Non-Value-Added Costs, Activity Volume and Unused Capacity Variances

Maquina Company produces custom-made machine parts. Maquina recently has implemented an activity-based management (ABM) system with the objective of reducing costs. Maquina has begun analyzing each activity to determine ways to increase its efficiency. Setting up equipment was among the first group of activities to be carefully studied. The study revealed that setup hours was a good driver for the activity. During the last year, the company incurred fixed setup costs of $588,800 (salaries of 16 employees). The fixed costs provide a capacity of 36,800 hours (2,300 per employee at practical capacity). The setup activity was viewed as necessary, and the value-added standard was set at 2,300 hours. Actual setup hours used in the most recent period were 34,700.

Required:

1. Calculate the volume and unused capacity variances for the setup activity. Enter all amounts as positive values.

Volume Variance $
  • Favorable
  • Unfavorable
Unused Capacity Variance $
  • Favorable
  • Unfavorable

2. Prepare a report that presents value-added, non-value-added, and actual costs for setup.

Maquina Company
Value- and Non-Value-Added Cost Report
Value-Added Non-Value-Added Actual
Setting up $ $ $
Answer #1
Part-1
Fixed Activity Rate (SP) =Fixed Setup Cost/ No. of Hour
$588800/36800=$16/Set upHour
Volume Variance= (SPXSQ-SPX AQ)
(16X2300)-(16X 36800)=$552000 UF
Unused Capacity Variance= (SPXAQ)-(SPX AU)
(16X36800)- (16X 34700)=33600F
Part-2: Value & Non-Value Added Cost Report
Value Added Non Value Added Actual
Setting Up (SPXSQ) Volume Variance SPX AQ
(16X 2300) (16X36800)
                      36,800                        5,52,000            5,88,800